# Ali Dasdan KD Consulting Saratoga, CA, USA

### v2

@@ 3.5 Extension: Condition Intensity and Tragedy Severity
  • Low intensity: Few actors (e.g., n=3n=3-5 firms in an oligopoly)
  • Medium intensity: Moderate number of actors (e.g., n=20n=20-50)
  • High intensity: Large number of actors (e.g., thousands of firms, hundreds of nations)
Impact on tragedy severity: More actors make coordination exponentially more difficult. Each actor’s individual contribution to the collective good diminishes as $n$ increases, while the temptation to defect remains constant. Additionally, heterogeneity among actors (different sizes, capabilities, preferences) further intensifies coordination difficulty.
Examples:
  • Low: OPEC coordination among  13 major oil producers—difficult but achievable
  • High: Climate change with 195+ nations plus countless firms—extremely difficult
  • Extreme: Antibiotics resistance with billions of patients, millions of prescribers across medical and agricultural sectors globally—coordination structurally impossible at individual decision level. Bank runs with millions of independent depositors making simultaneous withdrawal decisions during panic.
@@
  • Low intensity: Small negative externalities affecting few others
  • Medium intensity: Moderate externalities with delayed effects
  • High intensity: Large, immediate externalities affecting all actors severely
Impact on tragedy severity: Larger externalities create stronger feedback loops. When one actor’s defection severely harms others, it triggers more aggressive competitive responses, accelerating the race to the bottom. The speed and magnitude of externality transmission determines how rapidly the tragedy unfolds.
Examples:
  • Low: Overfishing in a large ocean—gradual depletion over decades
  • High: Bank runs—panic spreads within hours, complete collapse within days
  • High: AI capabilities races—competitor’s breakthrough immediately threatens all others
  • Extreme: Security dilemmas—states respond to perceived threats before they materialize, creating anticipatory arms races. Intelligence about potential enemy programs triggers immediate counter-programs based on worst-case assumptions, with no waiting period for actual capability development.
@@
  • Low intensity: Defection provides modest advantage; cooperation causes minor disadvantage
  • Medium intensity: Defection provides significant advantage; cooperation causes substantial losses
  • High intensity: Defection is existentially necessary; cooperation risks elimination (bankruptcy, conquest, extinction)
Impact on tragedy severity: The intensity of existential stakes determines how strongly dominance holds. When cooperation means death or bankruptcy, actors will defect regardless of collective costs. Lower existential stakes create space for partial coordination or voluntary restraint.
Examples:
  • Low: Hockey helmets—competitive disadvantage is marginal visibility/comfort
  • Medium: Productivity competition—firms face market share loss and competitive pressure, but not immediate bankruptcy. European firms reduced hours and survived, demonstrating stakes are substantial but not existential. Current AI competition exhibits similar dynamics: Google recovered from Transformer lag, multiple competitors remain viable despite capability gaps (ChatGPT, Gemini, DeepSeek).
  • High: Security dilemmas in anarchy—states face conquest and political death
  • Extreme: Transformative AI (AGI/superintelligence)— actors face permanent strategic obsolescence with no recovery possibility, or existential risk from misaligned systems. Unlike current competition where multiple players survive capability gaps, transformative breakthroughs may create winner-take-all dynamics where falling behind means permanent irrelevance.
@@
  • Low intensity: Universal cooperation slightly better than universal defection
  • Medium intensity: Universal cooperation moderately better—significant welfare gains available
  • High intensity: Universal cooperation vastly superior—catastrophic outcomes from mutual defection
Impact on tragedy severity: Greater Pareto-inefficiency means larger collective losses from coordination failure. This affects both the normative urgency (how important it is to solve) and practical difficulty (actors may invest more in solving severe tragedies). However, high inefficiency alone doesn’t make coordination easier—it only makes failure more costly.
Examples:
  • Low: Minor commons problems—modest resource depletion
  • Medium: Productivity competition—substantial leisure forgone, wage stagnation. Significant quality-of-life loss but society continues functioning with material prosperity.
  • High: Regional conflicts—major destruction but potential for recovery
  • Extreme: Climate change—tipping points create irreversible changes over millennia; ecosystem collapse and mass extinction cannot be undone on human timescales. Nuclear security dilemmas—potential human extinction from nuclear war. AGI misalignment—potential permanent loss of human control or extinction.
@@
  • Low intensity: Enforcement difficult but feasible; monitoring possible with effort and cost
  • Medium intensity: Enforcement very difficult; verification problematic; commitment fragile
  • High intensity: Enforcement structurally impossible; verification impossible; no credible commitment mechanism exists
Impact on tragedy severity: The impossibility of enforcement determines whether the tragedy is merely difficult or truly structural. When enforcement is difficult but possible (low intensity), sufficient institutional investment can achieve coordination. When enforcement is structurally impossible (high intensity), no amount of effort suffices without changing the underlying structure.
@@
Examples:
  • Low: National labor regulations—government can monitor and enforce within borders
  • Medium: Nuclear non-proliferation—verification possible but imperfect through inspections
  • High: International labor standards—no global authority, capital mobility enables evasion
  • Extreme: AI development—software is copyable, training is concealable, dual-use is inseparable
@@ 3.6 Implications of Intensity Analysis
Comparative prediction: We can predict relative difficulty of coordination across domains by comparing condition intensities. AI governance (for transformative breakthroughs) faces extreme intensity on three conditions simultaneously (C3, C4, C5), while severe historical cases like nuclear security dilemmas faced extreme intensity on two conditions (C2, C4) and climate change on one (C4). Current AI capabilities competition exhibits high but not extreme intensity, as evidenced by multiple viable competitors despite capability gaps.
Intervention prioritization: Interventions should target the highest-intensity conditions. For AI, the extreme intensity of C5 (enforcement impossibility due to verification) suggests compute governance (making verification partially possible) may be most promising despite remaining severe challenges.
Partial coordination explanation: The European case shows lower intensity on some conditions (reduced existential stakes through wealthy welfare states, supranational governance reducing anarchy) enables partial coordination despite high intensity on others (global competition, verification difficulty).
Temporal variation explanation: Condition intensities change over time. Productivity competition shows moderate base intensity (C2, C3 at medium due to gradual market dynamics and non-existential stakes), but globalization after 1980 increased N (more actors), strengthened externalities (faster competitive transmission), and weakened enforcement (capital mobility), making even this moderate tragedy harder to address. Climate change intensified as C4 shifted from high to extreme with growing evidence of irreversible tipping points.
Falsification refinement: The framework predicts that reducing condition intensity should reduce tragedy severity proportionally. Evidence of coordination success without reducing any condition’s intensity would falsify the framework. The European case confirms: partial reduction in intensity (some conditions) produces partial coordination (some dimensions).
@@ 3.7 Tragedy Index: Intensity Quantification
  • Structural Fingerprinting: As illustrated in Figure 1, plotting intensities on a radar chart allows us to visualize the ’shape’ of the coordination failure. This visualization reveals that AI Governance effectively “envelopes” historical cases like Nuclear Security. While Nuclear Security spikes on specific dimensions (Externalities and Pareto-Inefficiency), AI Governance maintains high or extreme intensity across all five dimensions, creating a unique and formidable structural profile.
    Figure 1: The Structural Fingerprint of AI Governance. Comparison of condition intensities for four selected domains (logarithmic scale), chosen to illustrate diverse structural profiles and highlight AI Governance’s unique fingerprint. Note how AI Governance (red) envelopes other tragedies, exhibiting “Extreme” intensity on three dimensions simultaneously (C3, C4, C5), whereas historical cases like Security or Climate Change spike only on specific dimensions. Overlapping areas (blended colors) indicate shared structural challenges.
@@
We propose the Tragedy Index ($T$) as a heuristic quantification of coordination difficulty. Two methodological choices deserve acknowledgment:
Exponential scaling: We use base-10 exponential intensities ($I_{c}\in\{0,10,10^{2},10^{3},10^{4}\}$) because coordination difficulty plausibly scales non-linearly: a tragedy with “Extreme” conditions on multiple dimensions may be orders of magnitude harder than one with “low” conditions, not merely several times harder. The base-10 choice is somewhat arbitrary but enables intuitive order-of-magnitude interpretation.
Geometric aggregation: We combine intensities via geometric mean (the 5th root of the product):
$T=\left(\prod_{c=1}^{5}I_{c}\right)^{\frac{1}{5}}$
where $I_{c}$ corresponds to Solved/Absent (0), Low (10), Medium (100), High (1,000), or Extreme (10,000). This multiplicative formulation handles the zero case naturally (solved tragedies yield $T=0$) and penalizes extreme conditions heavily, reflecting their disproportionate contribution to coordination difficulty. Alternative aggregations such as logarithmic sum would yield different absolute values but similar ordinal rankings.
@@
Under this heuristic measure, AI Governance (T=3,981T=3{,}981) faces orders-of-magnitude greater coordination difficulty than other severe tragedies. Climate change, bank runs, and nuclear security dilemmas all face catastrophic stakes (T=1, [^0] T=1{,}000 each), but differ in their condition profiles: security achieves this through anticipatory dynamics and extinction-level risks despite few actors; climate through irreversibility; bank runs through mass-scale coordination failure. Productivity competition ($T=251$) and antibiotics resistance ($T=398$), despite their global scale and persistence, face lower intensity due to non-existential stakes and gradual dynamics. Manageable cases like public goods ($T=40$) have been successfully addressed through institutional solutions. This quantification reinforces the qualitative analysis: AI governance presents an unprecedented coordination challenge.
Table 2 compares condition intensities across our main cases.
Table 2: Condition Intensity Comparison Across Domains, including the two new ones introduced in this paper. Intensities: L=Low (10), M=Medium (100), H=High (1,000), E=Extreme (10,000).a
Domain C1:N C2:B C3:D C4:P C5:E Index ($T$)
@@ 4 The Tragedy of Productivity
Although worker welfare includes wages and benefits in addition to working hours, we will use working hours as a proxy for worker welfare as well as to simplify the formal analysis.
@@ 4.2 Formalization: The N-Firm Game
    • $U_{F,i}(C,\mathbf{s_{-i}})=\pi\cdot\frac{Q_{i}}{Q_{i}+Q_{-i}(\mathbf{s_{-i}})}-c\cdot\frac{H_{i}}{\alpha}$
@@
    • $U_{F,i}(D,\mathbf{s_{-i}})=\pi\cdot\frac{\alpha Q_{i}}{\alpha Q_{i}+Q_{-i}(\mathbf{s_{-i}})}-c\cdot H_{i}$
@@
For any firm $i$, any competitor strategy profile $s_{-i}$ resulting in total competitor output $Q_{-i}(s_{-i})>0$, and any productivity factor $\alpha>1$:
@@
    Defection is the dominant strategy, i.e., $U_{i}(D,s_{-i})>U_{i}(C,s_{-i})$ for all $s_{-i}$, if the profit gain from this increased market share always exceeds the (constant) net cost of forgoing leisure.
@@
$\quad-\underbrace{\left(c\cdot H_{i}-c\cdot\frac{H_{i}}{\alpha}\right)}_{\text{Increased Labor Cost (Always Positive)}}$
@@
$U_{F,i}(D)-U_{F,i}(C)=\text{Market Share Gain}-\text{Increase Labor Cost}$
Defection is the dominant strategy (C3 is met) with the assumption that the profit from the Market Share Gain is greater than the Increased Labor Cost from not reducing hours.
@@ 2. For Workers ($U_{W}$):
Result: Combined with the N-player structure, externalities, and enforcement difficulty established in Section 4, productivity competition satisfies the complete set of conditions for structural tragedy. The coordination failure emerges from the fundamental competitive structure of markets facing productivity shocks, not from special assumptions about firm homogeneity or market structure.

### v3

@@ 3.5 Extension: Condition Intensity and Tragedy Severity
  • Low intensity: Few actors with stable identities (e.g., small groups like hockey leagues, bilateral security relationships)
  • Medium intensity: Moderate number of organized actors (e.g., dozens of states or firms in defined sectors like commons governance, public goods provision)
  • High intensity: Large number of heterogeneous actors across sectors (e.g., hundreds of nations plus countless firms in climate change, global productivity competition)
  • Extreme intensity: Massive number of individual-level decisions across decentralized populations (e.g., millions of depositors in bank runs, billions of patients and prescribers in antibiotics resistance)
Impact on tragedy severity:
More actors make coordination exponentially more difficult. Each actor’s individual contribution to the collective good diminishes as $n$ increases, while the temptation to defect remains constant. Additionally, heterogeneity among actors (different sizes, capabilities, preferences) further intensifies coordination difficulty.
@@
  • Low intensity: Small, localized negative externalities with delayed or uncertain effects (e.g., individual free-riding on public goods, minor competitive disadvantages in hockey helmets)
  • Medium intensity: Moderate externalities with gradual accumulation (e.g., commons depletion over years, antibiotic resistance developing gradually)
  • High intensity: Large, immediate externalities with rapid transmission (e.g., financial panic spreading within days, climate damages affecting all actors, AI capability advantages immediately threatening competitors)
  • Extreme intensity: Anticipatory externalities where actors respond to perceived threats before they materialize, creating self-reinforcing spirals (e.g., security dilemmas where intelligence about potential programs triggers immediate counter-programs)
Impact on tragedy severity:
Larger externalities create stronger feedback loops. When one actor’s defection severely harms others, it triggers more aggressive competitive responses, accelerating the race to the bottom. The speed and magnitude of externality transmission determines how rapidly the tragedy unfolds.
@@
  • Low intensity: Defection provides modest advantage; cooperation causes minor competitive disadvantage (e.g., visibility or comfort in hockey helmets, reduced returns in public goods)
  • Medium intensity: Defection provides significant competitive advantage; cooperation causes substantial market share loss but not immediate elimination (e.g., productivity competition where firms can survive with reduced hours, commons where individuals can survive with restraint, climate where nations face competitive disadvantage but can adapt)
  • High intensity: Defection necessary to avoid severe consequences; cooperation risks major losses (e.g., state conquest in security dilemmas, losing all deposits in bank runs)
  • Extreme intensity: Defection necessary to avoid permanent strategic obsolescence; restraint means irreversible loss with no possibility of recovery (e.g., transformative AI where falling behind creates permanent winner-take-all dynamics)
Impact on tragedy severity:
The intensity of existential stakes determines how strongly dominance holds. When cooperation means death or bankruptcy, actors will defect regardless of collective costs. Lower existential stakes create space for partial coordination or voluntary restraint.
@@
  • Low intensity: Universal cooperation slightly better than universal defection; modest welfare gains available (e.g., marginal safety improvements from universal helmet use)
  • Medium intensity: Universal cooperation moderately better; significant quality-of-life improvements but society continues functioning (e.g., commons preservation, public goods provision, productivity welfare gains, avoiding bank runs, maintaining antibiotic effectiveness)
  • High intensity: Universal cooperation vastly superior; defection leads to major catastrophic outcomes with potential for recovery (e.g., regional conflicts, economic depressions)
  • Extreme intensity: Universal cooperation necessary to avoid irreversible catastrophic outcomes with permanent losses (e.g., human extinction from nuclear war or AI misalignment, irreversible climate tipping points causing ecosystem collapse)
Impact on tragedy severity:
Greater Pareto-inefficiency means larger collective losses from coordination failure. This affects both the normative urgency (how important it is to solve) and practical difficulty (actors may invest more in solving severe tragedies). However, high inefficiency alone doesn’t make coordination easier; it only makes failure more costly.
@@
  • Low intensity: Enforcement absent or weak in current form but structurally feasible; clear authority could monitor and sanction violations with institutional development
  • Medium intensity: Enforcement difficult but possible with institutional investment; monitoring costly but achievable; commitment requires sustained effort (e.g., league rules for helmets, taxation for public goods)
  • High intensity: Enforcement very difficult due to international anarchy or distributed nature; verification problematic; commitment fragile; sanctions difficult but not impossible (e.g., commons without property rights, international labor standards, nuclear non-proliferation, climate agreements, bank run coordination)
  • Extreme intensity: Enforcement structurally impossible due to fundamental technical barriers (perfect hiding, instant copying, inseparable dual-use) combined with international anarchy and no feasible global authority (e.g., AI development where software is copyable, training concealable, capabilities unverifiable)
Impact on tragedy severity:
The impossibility of enforcement determines whether the tragedy is merely difficult or truly structural. When enforcement is difficult but possible (low to medium intensity), sufficient institutional investment can achieve coordination. When enforcement is structurally impossible (extreme intensity), no amount of effort suffices without changing the underlying structure.
@@ 3.6 Implications of Intensity Analysis
Comparative Prediction:
We can predict relative difficulty of coordination across domains by comparing condition intensities. AI governance (for transformative breakthroughs) faces extreme intensity on three conditions simultaneously (C3, C4, C5), while severe historical cases like nuclear security dilemmas faced extreme intensity on two conditions (C2, C4) and climate change on one (C4). Current AI capabilities competition exhibits high but not extreme intensity, as evidenced by multiple viable competitors despite capability gaps.
Intervention Prioritization:
Interventions should target the highest-intensity conditions. For AI, the extreme intensity of C5 (enforcement impossibility due to verification) suggests compute governance (making verification partially possible) may be most promising despite remaining severe challenges.
Partial Coordination Explanation:
The European case shows lower intensity on some conditions (reduced existential stakes through wealthy welfare states, supranational governance reducing anarchy) enables partial coordination despite high intensity on others (global competition, verification difficulty).
Temporal Variation Explanation:
Condition intensities change over time. Productivity competition shows moderate base intensity (C2, C3 at medium due to gradual market dynamics and non-existential stakes), but globalization after 1980 increased N (more actors), strengthened externalities (faster competitive transmission), and weakened enforcement (capital mobility), making even this moderate tragedy harder to address. Climate change intensified as C4 shifted from high to extreme with growing evidence of irreversible tipping points.
Falsification Refinement:
The framework predicts that reducing condition intensity should reduce tragedy severity proportionally. Evidence of coordination success without reducing any condition’s intensity would falsify the framework. The European case confirms: partial reduction in intensity (some conditions) produces partial coordination (some dimensions).
@@ 3.7 Tragedy Index: Intensity Quantification
  • Structural Fingerprinting: As illustrated in Figure 1, plotting intensities on a radar chart allows us to visualize the ‘shape’ of the coordination failure. This visualization reveals that AI Governance effectively ‘envelopes’ historical cases like Nuclear Security. While Nuclear Security spikes on specific dimensions (Externalities and Pareto-Inefficiency), AI Governance maintains high or extreme intensity across all five dimensions, creating a unique and formidable structural profile.
    Figure 1: The Structural Fingerprint of AI Governance. Comparison of condition intensities for four selected domains (logarithmic scale), chosen to illustrate diverse structural profiles and highlight AI Governance’s unique fingerprint. Note how AI Governance (red) envelopes other tragedies, exhibiting ‘Extreme’ intensity on three dimensions simultaneously (C3, C4, C5), whereas historical cases like Security or Climate Change spike only on specific dimensions. Overlapping areas (blended colors) indicate shared structural challenges.
@@
We propose the Tragedy Index ($\iota$) as a heuristic quantification of coordination difficulty. Two methodological choices deserve acknowledgment:
Exponential Scaling:
We use base-10 exponential intensities ($I_{c}\in\{0,10,10^{2},10^{3},10^{4}\}$) because coordination difficulty plausibly scales non-linearly: a tragedy with ‘Extreme’ conditions on multiple dimensions may be orders of magnitude harder than one with ‘low’ conditions, not merely several times harder. The base-10 choice is somewhat arbitrary but enables intuitive order-of-magnitude interpretation.
Geometric Aggregation:
We combine intensities via geometric mean (the 5th root of the product):
$\iota=\left(\prod_{c=1}^{5}I_{c}\right)^{\frac{1}{5}}$
where $I_{c}$ corresponds to Solved/Absent (0), Low (10), Medium (100), High (1,000), or Extreme (10,000).
We employ the geometric mean because the five conditions are necessary: the tragedy structure requires the conjoint presence of all factors. Mathematically, necessity implies logical conjunction, which corresponds to multiplication rather than addition. If any condition intensity falls to zero, the index correctly yields zero: no tragedy exists. An arithmetic mean would incorrectly assign positive severity when a necessary condition is absent.
Additionally, the nth-root operation penalizes extreme conditions heavily: a single extreme-intensity condition (10,000) contributes less to the index than moderate intensity across multiple conditions, reflecting the disproportionate coordination difficulty when several conditions bind simultaneously. Alternative aggregations such as logarithmic sum would yield different absolute values but preserve ordinal rankings across domains.
@@
As shown in Table 2, under this heuristic measure, AI Governance ($\iota=3{,}981$) faces orders-of-magnitude greater coordination difficulty than other severe tragedies. Climate change, bank runs, and nuclear security dilemmas all face catastrophic stakes ($\iota=1{,}000$ each), but differ in their condition profiles: security achieves this through anticipatory dynamics and extinction-level risks despite few actors; climate through irreversibility; bank runs through mass-scale coordination failure. Productivity competition ($\iota=251$) and antibiotics resistance ($\iota=398$), despite their global scale and persistence, face lower intensity due to non-existential stakes and gradual dynamics. Manageable cases like public goods ($\iota=40$) have been successfully addressed through institutional solutions. This quantification reinforces the qualitative analysis: AI governance presents an unprecedented coordination challenge.
Table 2: Intensity Comparison of Conditions C1-C5 Across Domains, including the two new ones introduced in this paper. Intensities: L=Low (10), M=Medium (100), H=High (1,000), E=Extreme (10,000).a
Domain C1 C2 C3 C4 C5 Index ($\iota$)
@@
3.8 Alternative Formalization: Game-Theoretic Measures
While the Tragedy Index provides intuitive comparison across domains through structural condition intensities, formal game-theoretic measures offer an alternative approach to quantifying tragedy severity. These measures focus on the payoff structure (Conditions 3 and 4) and provide mathematical precision about welfare losses and incentive misalignments.
Consider an N-player game satisfying Conditions 1-5. Let $U_{coop}$ denote global welfare under universal cooperation (profile $s^{C}$), $U_{nash}$ denote global welfare at the Nash equilibrium (profile $s^{D}$), and for symmetric games, let $T$ be the temptation payoff (defect while others cooperate), $R$ be the reward payoff (mutual cooperation), $P$ be the punishment payoff (mutual defection), and $S$ be the sucker payoff (cooperate while others defect). Structural tragedies satisfy the prisoner’s dilemma ordering: $T>R>P>S$.
The Normalized Rationality Gap:
(Nisan et al., 2007) measures the fraction of potential welfare lost due to coordination failure:
$\Delta_{norm}=\frac{U_{coop}-U_{nash}}{U_{coop}}=1-\frac{U_{nash}}{U_{coop}}$ (1)
This takes values in $[0,1]$, where $\Delta_{norm}\approx 0$ indicates minimal tragedy and $\Delta_{norm}\to 1$ indicates catastrophic welfare loss. For symmetric games, this simplifies to $\Delta_{norm}=(R-P)/R$.
The Price of Anarchy:
(Koutsoupias and Papadimitriou, 1999) measures the ratio between optimal social welfare and worst-case equilibrium welfare:
$\text{PoA}=\frac{U_{coop}}{U_{nash}}=\frac{R}{P}$ (2)
This multiplicative formulation naturally captures orders-of-magnitude differences: PoA $\approx$ 1 indicates near-efficiency, PoA $\in[10,100]$ indicates severe inefficiency, and PoA $\to\infty$ indicates existential tragedy where equilibrium approaches zero value.
The Temptation-Cooperation Ratio:
captures the tension between individual incentives and collective benefits (Rapoport and Chammah, 1965):
$\tau=\frac{T-R}{R-P}$ (3)
where $T$ is the temptation payoff (defect while others cooperate). This ratio measures how much individual gain from defection exceeds collective gain from cooperation: $\tau<1$ indicates coordination is structurally feasible, $\tau\approx 1$–5 indicates substantial coordination difficulty, and $\tau\to\infty$ indicates existential stakes make defection imperative.
These measures require detailed payoff specifications for each domain. For productivity competition, our formalization in Section 4 provides the foundation for such calculations. For AI governance, security dilemmas, and climate change, comprehensive payoff structures would require modeling uncertainty about catastrophic outcomes, discount rates, and long-term value functions—substantial undertakings beyond our current scope.
The Tragedy Index (Subsection 3.7) and game-theoretic measures serve complementary purposes: the Index enables systematic comparison across diverse domains through structural condition intensities without requiring precise payoff quantification, while game-theoretic measures provide rigorous welfare analysis when detailed payoff structures can be reliably estimated. Future research could apply these formal measures systematically across domains to validate and refine the Index’s ordinal rankings.
@@ 4 The Tragedy of Productivity
Although worker welfare includes wages and benefits in addition to working hours, we will use working hours as a proxy for worker welfare as well as to simplify the formal analysis. Our welfare analysis focuses on firms and workers as the direct participants in labor market coordination. Productivity gains may generate broader societal effects (e.g., consumer surplus from lower prices), but these lie outside our scope; we note that workers and consumers substantially overlap, making the welfare comparison largely internal to the same population.
@@ 4.2 Formalization: The N-Firm Game
    • $U_{F,i}(C,s_{-i})=\pi\cdot\frac{Q_{i}}{Q_{i}+Q_{-i}(s_{-i})}-c\cdot\frac{H_{i}}{\alpha}$
@@
    • $U_{F,i}(D,s_{-i})=\pi\cdot\frac{\alpha Q_{i}}{\alpha Q_{i}+Q_{-i}(s_{-i})}-c\cdot H_{i}$
@@
For any firm $i$, competitor strategy profile $s_{-i}$ with total competitor output $Q_{-i}(s_{-i})>0$, and productivity factor $\alpha>1$:
@@
    Defection is the dominant strategy, i.e., $U_{i}(D,s_{-i})>U_{i}(C,s_{-i})$ for all $s_{-i}$, provided the marginal profit from captured market share exceeds the marginal value of foregone leisure.
@@
$\quad-\underbrace{\left(c\cdot H_{i}-c\cdot\frac{H_{i}}{\alpha}\right)}_{\text{Labor Cost Increase (Always Positive)}}$
@@
$U_{F,i}(D)-U_{F,i}(C)=\text{Market Share Gain}-\text{Labor Cost Increase}$
Defection is the dominant strategy (C3 is met) with the assumption that the profit from the Market Share Gain is greater than the Labor Cost Increase from not reducing hours.
Assumption Justification:
@@
Note on Market Profit:
The proof assumes total market profit $\pi$ remains constant for analytical clarity. In practice, universal output expansion would likely reduce industry profit through price competition, meaning the actual welfare loss from mutual defection exceeds our estimate. Our formalization therefore provides a conservative lower bound on the tragedy’s severity.
Result — Structural Tragedy Proven:
Combined with the N-player structure, externalities, and enforcement difficulty established in Section 4, productivity competition satisfies the complete set of conditions for structural tragedy. The coordination failure emerges from the fundamental competitive structure of markets facing productivity shocks, not from special assumptions about firm homogeneity or market structure.